OpenStudy (anonymous):

Gabby solved x^2 - 3x = 4 as shown below: x^2 - 3x - 4 = 0 (x - 4)(x + 1) = 0 x = -4 and x = 1 Caitlyn says Gabby is wrong and believes this is the solution: x2 - 3x - 4 = 0 (x - 4)(x + 1) = 0 x = 4 and x = -1 Is Gabby or Caitlyn correct? What mistake was made by the incorrect girl?

OpenStudy (johnweldon1993):

Caitlyn is correct here, When you are down to \[\large (x - 4)(x + 1)=0\] you find out when that equation equals 0 Since 2 thing are being multiplied, we know that is either one is 0...anything multiplied by it will still equate to 0 So that means..either (x - 4) will equal 0 or (x + 1) will equal 0 When does that happen in both cases? Well (x - 4) = 0 happens when x = 4 and (x + 1) = 0 when x = -1